### v1

@@ 2.3 Artifact-reduced 3D reconstruction network
Three-dimensional reconstruction is another significant application of ETNet. However, direct application of ET2dNet to WF image stacks generates artifacts characterized by intensity discontinuities along the z-axis (see Supplementary Information Fig. S3). To reduce these artifacts, we employed knowledge distillation [wang_knowledge_2022] to train ET3dNet, a dedicated 3D reconstruction network, by transferring knowledge from ET2dNet.
@@
ET3dNet was trained on 881 EPI image stacks (6 layers each) using the hybrid architecture depicted in Fig. 3a. Given the absence of 3D GT, we generated pseudo-GT volumes using ET2dNet predictions from individual EPI layers. ET3dNet simultaneously outputs a reconstructed fake GT stack and a background stack. Supervised loss was computed between the reconstructed fake GT and ET2dNet pseudo-GT. The reconstructed fake GT was convolved with the PSF and combined with the background stack to generate fake EPI stacks, enabling self-supervised loss calculation against real EPI stacks. The network backbone utilized EViT-Unet [li_evit-unet_2025]. For more details see Methods 4.4.
@@
Figure 3: Characterizations and demonstrations of ET3dNet. a The hybrid architecture of ET3dNet and the knowledge distillation process. b Representative xy (top) and xz (bottom) slice of F-actin labeled C6 cell stack, acquired by EPI(first column), DW 3D reconstruction(second column), Self-Net (third column), UniFMIR(fourth column), ET3dNet (fifth column), and Sparse Deconvolution reconstruction of ET3dNet output (sixth column). c Three-color lateral maximum-intensity projections of a fixed C6 cell, maximum-intensity projections (top) and representative xz slice (bottom), acquired by EPI(first column), DW 3D reconstruction(second column), Self-Net (third column), UniFMIR(fourth column), ET3dNet (fifth column), and Sparse Deconvolution reconstruction of ET3dNet output (sixth column). Red, F-actin labeled with Alexa Fluor™ 635 Phalloidin; blue, membrane labeled with Alexa Fluor™ 488 WGA; purple, nucleus stained by Hoechst 33258. d Representative xy slice of human brain tissue stack containing 6-layers, acquired by EPI(first column), DW (second column), confocal(fourth column), ET3dNet (fifth column), and Sparse Deconvolution reconstruction of ET3dNet output (sixth column). And DW of stacks containing 41-layers (third column). Scale bar: 10 $\rm\mu m$.
2.3.2 comparison of ET3dNet and SOTA 3D reconstruction algorithm
@@
We assessed the performance of ET3dNet, and made comparisons to SOTA 3D reconstruction algorithms (Fig.3 b, c, d), including self-supervised learning method Self-Net [ning_deep_2023], supervised learning method UniFMIR [ma_pretraining_2024] and volumetric deconvolution method DW[wernersson_deconwolf_2024]. Implementation details of DW, Self-Net and UniFMIR are provided in Supplementary Information Note 3. Implementations on F-actin-labeled C6 cells ( Fig.3 b) revealed that neither UniFMIR nor Self-Net achieved isotropic resolution. We supposed that UniFMIR’s performance suffered from unrealistic synthetic training data. Self-Net requires clear lateral images, while the strong out-of-focus in such densely labeled sample deteriorates its performance. Although DW deconvolution outperformed learning-based methods, residual out-of-focus signal persisted in xz-planes (Fig. 3b). In contrast, ET3dNet produced images with substantially reduced background and out-of-focus signals across both xy- and xz-planes. Notably, the ET3dNet reconstruction exhibited lower lateral sharpness than the DW outputs. However, this limitation could be mitigated by integrating ET3dNet with Sparse Deconvolution. For implementation details, refer to Supplementary Information Note 2.
@@
In the comparative analysis on triple-labeled C6 cells (F-actin, membrane, nucleus; Fig. 3c), the ET3dNet reconstructions preserved finer membrane structural details, confirming its superiority over learning-based methods. In xz-planes, ET3dNet outputs revealed distinct arc-shaped membrane curvatures, indicating significantly improved defocus suppression. Furthermore, ET3dNet demonstrated effective compatibility with Sparse Deconvolution. Collectively, ET3dNet outperformed SOTA learning-based techniques and was comparable to SOTA deconvolution method for axial-resolution enhancement capability.
Real-time WF tomographic imaging faces a fundamental challenge in 3D reconstruction, where the requirement for numerous z-slices to achieve faithful reconstruction inherently compromises imaging speed. Particularly, in 3D reconstruction, background suppression is most effective when the defocused background from each slice is treated as part of the entire reconstructed volume rather than being removed in isolation (see Supplementary Information Note 4). Moreover, long exposure time can also cause photobleaching. Therefore, balancing reconstruction fidelity and acquisition scale is critical for practical biological applications. ET3dNet alleviated this by significantly reducing the slice requirement to only 6 layers, as shown in Fig. 3d. Under such limited-sampling conditions, DW deconvolution failed to adequately identify out-of-focus signals, as illustrated by the comparison of the DW results of 6-layer stack and 41-layer stack in Fig. 3d. This reduced slice dependency facilitated fast and accurate 3D reconstruction, making ET3dNet particularly advantageous for live-cell imaging.
@@ 3 Discussion
Our study established a deep learning-enabled EPI-TIRF cross-modality paradigm that computationally translates conventional WF images into TIRF-quality outputs without any hardware modification. ET2dNet overcomes TIRF’s shallow penetration limit, while achieving TIRF-comparable background suppression and axial super-resolution from single 2D WF images. The hybrid architecture of ET2dNet synergizes supervised learning (using real EPI-TIRF pairs) with physics-based self-supervision (via convolution with the PSF), and can improve generalization and facilitate finetuning. For 3D reconstruction, ET3dNet extends this capability through knowledge distillation, generating artifact-reduced volumes from z-stacks.
The axial resolution of our current method can still be further improved. Here, the capability is fundamentally constrained by the quality of the TIRF images in our training dataset. Advances in TIRF-based technology, such as multi-angle [fu_axial_2016] and two-wavelength [stabley_real-time_2015] TIRF, can achieve nearly 20nm axial resolution. These methods could be incorporated into future versions. To quantify the ultimate resolution in such a case, our current validation, which relies on indirect biological evidence (e.g., arch-shaped microfilament reconstruction) and comparative analysis with TIRF/confocal references is insufficient. More rigorous and quantitative evaluation is required. This could be achieved by adapting lateral resolution quantification approaches (e.g., Fourier Ring Correlation analysis [koho_fourier_2019, zhao_quantitatively_2023] or Decorrelation method [descloux_parameter-free_2019]) to the axial dimension via some specific modification. Beyond improving the training dataset, employing some effective alternative network architectures, such as RCAN [chen_three-dimensional_2021], DFCAN [qiao_evaluation_2021], and GAN [wang_deep_2019], may further improve the performance of ETNet.
@@
Our work also contributes new insights to algorithm‐based super‐resolution. It is well recognized that task‐specific deep learning models often suffer from limited generalizability when applied to different restoration tasks [ma_pretraining_2024, moor_foundation_2023]. One promising strategy to address this is pretraining networks to learn priors across diverse tasks, which can then benefit new restoration goals [ma_pretraining_2024, moor_foundation_2023, liu_deep_2020]. In the cases of axial resolution‐enhanced restoration and related tasks, the EPI-TIRF cross-modality task serves as an excellent source of such priors. We demonstrated that these learned priors could be successfully transferred to yield improved performance, as evidenced by the knowledge distillation from ET2dNet to ET3dNet. This advantage could be extended to related applications, such as time-lapse super-resolution [qiao_neural_2025], through a similar knowledge distillation process. Collectively, we believe our work will facilitate biological discovery while promoting novel developments in computational super-resolution algorithms.
@@ 4.3.2 Training dataset enhancement and normalization
Due to the differences in illumination intensity and uniformity between the EPI and TIRF modes, the absolute intensity of EPI-TIRF images was not comparable. Hence, normalizations were performed on each EPI and TIRF image, where the intensity for each pixel in an image was divided by a maximum value to be in [0,1]. In practice, the percentile normalization was performed on the training dataset, which is formulated as:
$norm(\boldsymbol{I},p_{low},p_{high})=\frac{\boldsymbol{I}-perc(\boldsymbol{I},p_{low})}{perc(\boldsymbol{I},p_{high})-perc(\boldsymbol{I},p_{low})},$ (1)
where $perc(\boldsymbol{I},p)$ denotes the p-th percentile of the input image $\boldsymbol{I}$. $p_{low}$ and $p_{high}$ are typically set to 0 and 98, respectively, during training and testing. A variance threshold-based analysis was conducted to ensure that the EPI-TIRF image pairs contained sufficient information. Note that out-of-focus images, which may appear during inference of an image stack, were not included in the training dataset, and this may result in unexpected outputs. It is evident that the variance of an in-focus image is greater than that of an out-of-focus image. The problem of out-of-focus images was alleviated by multiplying each TIRF image by its own variance.
@@ 4.4 ETNet
Given an EPI-TIRF image pair, denoted as $\boldsymbol{x}$ (EPI) and $\boldsymbol{y}$ (TIRF), the network output layer was designed to have twice the number of channels as the input, enabling separate reconstruction of the ground truth (GT) and background components. Specifically, $(\hat{\boldsymbol{g}},\hat{\boldsymbol{b}})=f_{\theta}(\boldsymbol{x})$, where $f_{\theta}(\cdot)$ is the trainable model, $\hat{\boldsymbol{g}}$ is the predicted GT, and $\hat{\boldsymbol{b}}$ is the predicted background. Here and throughout, the hat symbol $\hat{}$ denotes network outputs.
For a single-layer EPI image $\boldsymbol{x}$, the output is a two-channel map. The reconstructed TIRF image is given by $\hat{\boldsymbol{y}}=\hat{\boldsymbol{g}}*psf$, and the reconstructed EPI image is obtained as $\hat{\boldsymbol{x}}=\hat{\boldsymbol{y}}+\hat{\boldsymbol{b}}$.
@@
where $\mathcal{L}(\cdot,\cdot)$ denotes the loss between two images, $\boldsymbol{y}$ is the TIRF image, $\boldsymbol{x}$ is the input EPI image, $\hat{\boldsymbol{y}}$ is the predicted TIRF image, and $\hat{\boldsymbol{x}}$ is the reconstructed EPI image. Here, $\mathcal{L}(\hat{\boldsymbol{y}},\boldsymbol{y})$ serves as the supervised loss, $\mathcal{L}(\hat{\boldsymbol{x}},\boldsymbol{x})$ acts as the self-supervised loss, and $\mathcal{R}$ is the regularization term. Each loss term is formulated as a weighted sum of several components, including L1 loss, L2 loss, total variation (TV) loss, and multi-scale structural similarity (MS-SSIM[Wang_2003]) loss:
@@
For ET3dNet, let the input EPI stack be denoted as $\boldsymbol{x}_{3d}$ (where the subscript $3d$ indicates a 3D stack), the predicted ground truth stack as $\hat{\boldsymbol{g}}_{3d}$, and the predicted background stack as $\hat{\boldsymbol{b}}_{3d}$. A 3D point spread function $psf_{3d}$ is used for convolution, resulting in the reconstructed TIRF stack: $\hat{\boldsymbol{y}}_{3d}=\hat{\boldsymbol{g}}_{3d}*psf_{3d}$. The reconstructed EPI stack is then given by $\hat{\boldsymbol{x}}_{3d}=\hat{\boldsymbol{y}}_{3d}+\hat{\boldsymbol{b}}_{3d}$.
Since true GT stacks are unavailable, we leveraged the pre-trained ET2dNet (with fixed parameters) to process each slice of the 3D EPI stack, generating a pseudo-GT stack $\boldsymbol{g}_{3d}^{\dagger}$. This pseudo-GT serves as the supervisory target for ET3dNet training. For details on training a teacher (ET2dNet) for ET3dNet, refer to Supplementary Information Note 6.
@@
where $\mathcal{L}(\hat{\boldsymbol{g}}_{3d},\boldsymbol{g}_{3d}^{\dagger})$ is the supervised loss between the predicted and pseudo-GT stacks, $\mathcal{L}(\hat{\boldsymbol{x}}_{3d},\boldsymbol{x}_{3d})$ is the self-supervised loss between the reconstructed and input EPI stacks, and $\mathcal{R}$ is the regularization term:

### v2

@@ 2.3 Volumetric computational TIRF reconstruction
While single-frame computational TIRF provides effective optical sectioning for focal planes, its extension to volumetric imaging introduces additional challenges along the axial dimension. In practice, volumetric reconstruction is typically performed from a finite stack, which can contain strong out-of-focus background from emitters located outside the reconstructed volume. In this setting, independently reconstructing each axial plane fails to enforce axial consistency and cannot jointly account for out-of-volume background that couples multiple planes. As a result, volumetric reconstructions can degrade in the presence of strong out-of-focus background (see examples Supplementary Information Fig. S3). To address these challenges, we extend cTIRF to volumetric reconstruction by leveraging axial context across the stack, and implement this extension using a dedicated three-dimensional network (ET3dNet) trained via knowledge distillation[wang_knowledge_2022] from ET2dNet .
@@
ET3dNet was trained on 881 EPI image stacks (6 layers each) using the hybrid architecture depicted in Fig. 3a. Given the absence of 3D GT, we generated slice-wise pseudo-GT volumes using ET2dNet predictions from individual EPI layers. ET3dNet simultaneously outputs a volumetric pseudo-GT stack and a background stack. Supervised loss was computed between the volumetric pseudo-GT and slice-wise pseudo-GT. The volumetric pseudo-GT was convolved with the PSF (generates cTIRF stack) and combined with the background stack to generate reconstructed EPI stacks, enabling self-supervised loss calculation against experimental measured EPI stacks. The network backbone utilized EViT-Unet [li_evit-unet_2025]. More details are shown in Methods 4.4.
@@
Figure 3: Characterizations and demonstrations of volumetric computational TIRF. a The hybrid architecture of ET3dNet and the knowledge distillation process. b Representative xy (top) and xz (bottom) slice of F-actin labeled C6 cell stack, acquired by EPI(first column), DW 3D reconstruction(second column), Self-Net (third column), UniFMIR(fourth column), volumetric cTIRF (ET3dNet outputs, fifth column), and Sparse Deconvolution reconstruction of volumetric cTIRF (sixth column). c Three-color lateral maximum-intensity projections of a fixed C6 cell, maximum-intensity projections (top) and representative xz slice (bottom), acquired by EPI(first column), DW 3D reconstruction(second column), Self-Net (third column), UniFMIR(fourth column), volumetric cTIRF (fifth column), and Sparse Deconvolution reconstruction of volumetric cTIRF (sixth column). Red, F-actin labeled with Alexa Fluor™ 635 Phalloidin; blue, membrane labeled with Alexa Fluor™ 488 WGA; purple, nucleus stained by Hoechst 33258. d Representative xy slice of human brain tissue stack containing 6-layers, acquired by EPI(first column), DW (second column), confocal(fourth column), volumetric cTIRF (fifth column), and Sparse Deconvolution reconstruction of volumetric cTIRF (sixth column). And DW of stacks containing 41-layers (third column). Scale bar: 10 $\rm\mu m$.
2.3.2 comparison of volumetric cTIRF and 3D reconstruction algorithm
@@
We assessed the performance of volumetric cTIRF, and made comparisons to SOTA 3D reconstruction algorithms (Fig.3 b, c, d), including self-supervised learning method Self-Net [ning_deep_2023], supervised learning method UniFMIR [ma_pretraining_2024] and volumetric deconvolution method DW[wernersson_deconwolf_2024]. Implementation details of DW, Self-Net and UniFMIR are provided in Supplementary Information Note 3. Implementations on F-actin-labeled C6 cells ( Fig.3 b) revealed that neither UniFMIR nor Self-Net achieved isotropic resolution. We supposed that UniFMIR’s performance suffered from unrealistic synthetic training data. Self-Net requires clear lateral images, while the strong out-of-focus in such densely labeled sample deteriorates its performance. Although DW deconvolution outperformed learning-based methods, residual out-of-focus signal persisted in xz-planes (Fig. 3b). In contrast, volumetric cTIRF produced images with substantially reduced background and out-of-focus signals across both xy- and xz-planes. Notably, the volumetric cTIRF reconstruction exhibited lower lateral sharpness than the DW outputs. However, this limitation could be mitigated by integrating volumetric cTIRF with Sparse Deconvolution. For implementation details, refer to Supplementary Information Note 2.
@@
In the comparative analysis on triple-labeled C6 cells (F-actin, membrane, nucleus; Fig. 3c), the volumetric cTIRF reconstructions preserved finer membrane structural details, confirming its superiority over learning-based methods. In xz-planes, volumetric cTIRF revealed distinct arc-shaped membrane curvatures, indicating significantly improved defocus suppression. Furthermore, volumetric cTIRF demonstrated effective compatibility with Sparse Deconvolution. Collectively, volumetric cTIRF outperformed SOTA learning-based techniques and was comparable to SOTA deconvolution method for axial-resolution enhancement capability.
Deconvolution-based volumetric reconstruction assumes that the recorded signal in a finite axial stack arises from emitters within the reconstructed volume. This assumption necessitates sufficient axial sampling to achieve accurate reconstruction (see Supplementary Information Note 4). This requirement, however, limits imaging speed and increases photobleaching. Volumetric cTIRF alleviates this constraint by enabling robust reconstruction from a limited sampled axial stack containing only six planes (Fig.3d). Under such limited-sampling conditions, DW deconvolution failed to remove out-of-focus background, as evidenced by the marked degradation observed when comparing reconstructions from 6-layer and 41-layer stacks(Fig.3d second and third columns). By reducing the dependency on sufficient axial sampling while maintaining effective background suppression, volumetric cTIRF enables fast and reliable three-dimensional imaging, making it particularly well suited for live-cell applications.
@@ 3 Discussion
In this study, we establish computational TIRF as a computational imaging modality that transforms widefield epifluorescence microscopy into a TIRF-like representations without hardware modification. The single-frame cTIRF overcomes the intrinsic depth limitation of hardware-based TIRF while achieving effective background suppression and axial super-resolution from single widefield images. The practical deployability of cTIRF is enabled by a hybrid learning strategy that integrates supervision from experimentally measured EPI–TIRF pairs with physics-informed self-consistency based on the optical PSF. This combination improves generalization across imaging conditions and facilitates rapid adaptation to new imaging systems with minimal additional data. Beyond single-frame imaging, cTIRF is further extended to volumetric reconstruction through a three-dimensional network trained via knowledge distillation, enabling artifact-reduced 3d reconstructions from limited sampled axial stacks.
The axial resolution of our current method can still be further improved. Here, the capability is fundamentally constrained by the quality of the TIRF images in our training dataset. Advances in TIRF-based technology, such as multi-angle [fu_axial_2016] and two-wavelength [stabley_real-time_2015] TIRF, can achieve nearly 20nm axial resolution. These methods could be incorporated into future versions. Beyond improving the training dataset, employing some effective alternative network architectures, such as RCAN [chen_three-dimensional_2021], DFCAN [qiao_evaluation_2021], and GAN [wang_deep_2019], may further improve the performance of cTIRF.
@@
Our work also contributes new insights to algorithm‐based super‐resolution. It is well recognized that task‐specific deep learning models often suffer from limited generalizability when applied to different restoration tasks [ma_pretraining_2024, moor_foundation_2023]. One promising strategy to address this is pretraining networks to learn priors across diverse tasks, which can then benefit new restoration goals [ma_pretraining_2024, moor_foundation_2023, liu_deep_2020]. In the cases of axial resolution‐enhanced restoration and related tasks, the EPI-TIRF cross-modality task serves as an excellent source of such priors. We demonstrated that these learned priors could be successfully transferred to yield improved performance, as evidenced by the knowledge distillation from ET2dNet to ET3dNet. This advantage could be extended to related applications, such as time-lapse super-resolution [qiao_neural_2025], through a similar knowledge distillation process that considering temporal consistency. Collectively, we believe our work will facilitate biological discovery while promoting novel developments in computational super-resolution algorithms.
@@ 4.3.2 Training dataset enhancement and normalization
Due to the differences in illumination intensity and uniformity between the EPI and TIRF modes, the absolute intensity of EPI-TIRF images was not comparable. Hence, normalizations were performed on each EPI and TIRF image, where the intensity for each pixel in an image was divided by a maximum value to be in [0,1]. In practice, the percentile normalization with clipping was performed on the training dataset, which is formulated as:
$norm(\boldsymbol{I},p_{high})=\min(1,\frac{\boldsymbol{I}}{perc(\boldsymbol{I},p_{high})})$ (1)
where $perc(\boldsymbol{I},p)$ denotes the p-th percentile of the input image $\boldsymbol{I}$. $p_{high}$ is typically set to 98 during training and testing. A variance threshold-based analysis was conducted to ensure that the EPI-TIRF image pairs contained sufficient information. Note that out-of-focus images, which may appear during inference of an image stack, were not included in the training dataset, and this may result in unexpected outputs. It is evident that the variance of an in-focus image is greater than that of an out-of-focus image. The problem of out-of-focus images was alleviated by multiplying each TIRF image by its own variance.
@@ 4.4 ETNet
Given an EPI-TIRF image pair, denoted as $\boldsymbol{x}$ (EPI) and $\boldsymbol{y}$ (TIRF), the network output layer was designed to have twice the number of channels as the input, enabling separate reconstruction of the ground truth (GT) and background components. Specifically, $(\hat{\boldsymbol{g}},\hat{\boldsymbol{b}})=f_{\theta}(\boldsymbol{x})$, where $f_{\theta}(\cdot)$ is the trainable model, $\hat{\boldsymbol{g}}$ is the pseudo-GT, and $\hat{\boldsymbol{b}}$ is the predicted background. Here and throughout, the hat symbol $\hat{}$ denotes network outputs.
For a single-layer EPI image $\boldsymbol{x}$, the output is a two-channel map. The cTIRF image is given by $\hat{\boldsymbol{y}}=\hat{\boldsymbol{g}}*psf$, and the reconstructed EPI image is obtained as $\hat{\boldsymbol{x}}=\hat{\boldsymbol{y}}+\hat{\boldsymbol{b}}$.
@@
where $\mathcal{L}(\cdot,\cdot)$ denotes the loss between two images, $\boldsymbol{y}$ is the TIRF image, $\boldsymbol{x}$ is the input EPI image, $\hat{\boldsymbol{y}}$ is the cTIRF image, and $\hat{\boldsymbol{x}}$ is the reconstructed EPI image. Here, $\mathcal{L}(\hat{\boldsymbol{y}},\boldsymbol{y})$ serves as the supervised loss, $\mathcal{L}(\hat{\boldsymbol{x}},\boldsymbol{x})$ acts as the self-supervised loss, and $\mathcal{R}$ is the regularization term. Each loss term is formulated as a weighted sum of several components, including L1 loss, L2 loss, total variation (TV) loss, and multi-scale structural similarity (MS-SSIM[Wang_2003]) loss:
@@
For ET3dNet, let the input EPI stack be denoted as $\boldsymbol{x}_{3d}$ (where the subscript $3d$ indicates a 3D stack), the predicted volumetric pseudo-GT stack as $\hat{\boldsymbol{g}}_{3d}$, and the predicted background stack as $\hat{\boldsymbol{b}}_{3d}$. A 3D point spread function $psf_{3d}$ is used for convolution, resulting in the cTIRF stack: $\hat{\boldsymbol{y}}_{3d}=\hat{\boldsymbol{g}}_{3d}*psf_{3d}$. The reconstructed EPI stack is then given by $\hat{\boldsymbol{x}}_{3d}=\hat{\boldsymbol{y}}_{3d}+\hat{\boldsymbol{b}}_{3d}$.
Since true GT stacks are unavailable, we leveraged the pre-trained ET2dNet (with fixed parameters) to process each slice of the 3D EPI stack, generating a slice-wise pseudo-GT stack $\boldsymbol{g}_{3d}^{\dagger}$. This slice-wise pseudo-GT serves as the supervisory target for ET3dNet training. For details on training a teacher (ET2dNet) for ET3dNet, refer to Supplementary Information Note 6.
@@
where $\mathcal{L}(\hat{\boldsymbol{g}}_{3d},\boldsymbol{g}_{3d}^{\dagger})$ denotes the supervised loss computed over the pseudo GT stacks generated by the teacher and student networks, $\mathcal{L}(\hat{\boldsymbol{x}}_{3d},\boldsymbol{x}_{3d})$ is the self-supervised loss between the reconstructed and input EPI stacks, and $\mathcal{R}$ is the regularization term:
@@ 4.6 PSF usage and generation
4.7 Axial resolution estimation using decorrelation analysis
Axial resolution was quantified using decorrelation analysis [descloux_parameter-free_2019], a single-image resolution estimator that determines the highest spatial frequency $k_{c}$ containing reliable signal and relates it to the effective resolution. Compared with other quantitative resolution assessment methods, such as Fourier ring correlation (FRC) analysis[koho_fourier_2019, zhao_quantitatively_2023], decorrelation analysis circumvents the need for an empirical threshold and can therefore be applied universally across different imaging modalities.
For the analysis, an EPI stack of 201 slices was acquired with an axial sampling of 100 nm. Each XY plane comprised 512×512 pixels with a lateral pixel size of 97.5 nm. ET2dNet inference was applied independently to each XY slice to generate a corresponding cTIRF stack. Prior to inference, pixel intensities for each XY slice were normalized by dividing by $perc(\boldsymbol{I},p_{high})$ (where $p_{high}=98$, see Eq.1), and were rescaled by the same factor after inference.
The raw EPI and cTIRF stacks were then re-sliced into 512 XZ and 512 YZ cross-sections (each 201×512 pixels). Sectorial resolution estimation was performed on each cross-section using eight angular sectors, and the cutoff frequency $k_{c}$ from the fifth sector, corresponding to the axial direction, was extracted. The mean $k_{c}$ values for the raw EPI stack were 0.19 (XZ) and 0.21 (YZ), whereas the corresponding values for the cTIRF stack were 0.60 (XZ) and 0.67 (YZ). These measurements correspond to an approximately threefold improvement in axial resolution. Further details are provided in Supplementary Information Note 7.